# A Safe Preference Learning Approach for Personalization with Applications to Autonomous Vehicles

### 22 Jan 2024 v2

@@ Abstract
This work introduces a preference learning method that ensures adherence to given specifications, with an application to autonomous vehicles. Our approach incorporates the priority ordering of Signal Temporal Logic (STL) formulas describing traffic rules into a learning framework. By leveraging Parametric Weighted Signal Temporal Logic (PWSTL), we formulate the problem of safety-guaranteed preference learning based on pairwise comparisons and propose an approach to solve this learning problem. Our approach finds a feasible valuation for the weights of the given PWSTL formula such that, with these weights, preferred signals have weighted quantitative satisfaction measures greater than their non-preferred counterparts. The feasible valuation of weights given by our approach leads to a weighted STL formula that can be used in correct-and-custom-by-construction controller synthesis. We demonstrate the performance of our method with a pilot human subject study in two different simulated driving scenarios involving a stop sign and a pedestrian crossing. Our approach yields competitive results compared to existing preference learning methods in terms of capturing preferences, and notably outperforms them when safety is considered.
@@ I Introduction
Preferences are a fundamental aspect of human behavior and decision-making, and it is valuable to design autonomous systems that allow for personalization to better suit the needs and desires of users. In particular for autonomous driving, surveys have demonstrated that drivers have different comfort and performance preferences while driving in different scenarios and conditions [2, 3, 4]. Moreover, drivers tend to prefer different driving styles for autonomous vehicles than their own styles [5]. Customizing autonomous vehicles based on user preferences can increase user satisfaction with these vehicles. However, autonomous systems often require satisfaction of a set of rules for safe operation. Relying on human preferences alone may result in unsafe behaviors. For instance, at an intersection with a stop sign, drivers may sometimes prefer a rolling stop, which is illegal, over a full stop. However, an autonomous vehicle should always stop completely at a stop sign to guarantee the safety of all agents in the environment. Preference learning algorithms for safety critical operations must consider rule satisfaction. The main motivation for our work is the need for safe, trustworthy and customizable autonomous vehicle algorithms.
For safety-critical applications like driving, there are three desirable properties a preference learning method should satisfy to allow safe personalization: (i) expressivity: the model should be expressive enough to capture preferences; (ii) safety: it ensures safety by preferring a rule-following behavior against a rule-violating one (even in cases where the latter is scarce in the training data); and (iii) usability in control design: the learned model should be easy to integrate into downstream correct-by-construction control synthesis tasks. In this work, we propose to incorporate personalization and safety in one framework by using Signal Temporal Logic (STL) in a way to satisfy all of these properties. STL is a variant of temporal logic that is tailored for reasoning about the temporal properties of time series data and commonly used in describing correct behaviors in autonomous systems. Its descriptive power on systems opens wide application areas [6, 7, 8, 9, 10, 11].
@@
To evaluate the performance of our framework, we simulate two different driving scenarios: one with an autonomous vehicle navigating an intersection with a stop sign and one with an autonomous vehicle approaching a crosswalk while there is a pedestrian crossing the road. We generate two sets of trajectories that comply with traffic rules for these scenarios, and run a pilot human subject study with eight participants for both scenarios. Comparisons with baseline preference learning methods verify the need for safety-aware preference learning by showing that baseline methods usually lead to unsafe selections, whereas our method does not.
@@ II Literature Review
On the other hand, encoding safety rules in temporal logic is an eminent method for safety-critical applications [20, 21]. Specifications in temporal logic can be used for controller synthesis [6, 7, 21], motion planning [8, 22, 23] and learning applications [24, 9, 10, 25, 26, 11, 27, 28] in many autonomous systems. In particular, works in [24, 9, 10, 25, 29] try to infer a temporal logic formula for classification from the data. As a subset of learning applications, in robot learning, Chou et al. [26] tries to learn task specifications in linear temporal logic from demonstrations, Puranic et al. [11] scores demonstrations with the help of ordered specifications in the form of signal temporal logic, and works in [27, 28] use temporal logic for reward shaping and reinforcement learning.
Incorporating preferences and priorities with temporal logics is studied in [12, 30, 31, 32, 1]. The work in [12] introduces a weighted variant of the STL, called Weighted Signal Temporal Logic (WSTL), in which weights reflect the order of priorities or preferences. The work in [30] defines Weighted Truncated Linear Temporal Logic. Both works assumes that they have the knowledge for the formula and associated weights. For the end-user, it is hard to interpret the weights and define their preferences in the temporal logic formalism, so there needs to be an intermediate step to infer the weights from the user. In [31, 32], a parametric extension of WSTL, which we call PWSTL, is used in a time series classification problem, where weights of the formula are learned using neural networks.
@@ III-A Signal Temporal Logic (STL)
STL is a temporal logic formalism used to reason about signals $s:\mathbb{T}\to\mathcal{S}$, where $\mathbb{T}$ is a time domain and $\mathcal{S}\subseteq{\mathbb{R}^{m}_{e}}$ is a $m$ dimensional extended real-valued signal domain [33]. We will consider $\mathbb{T}$ to be infinite $\mathbb{Z}_{\geq 0}$ or finite $[0,t_{final}]\subset\mathbb{Z}_{\geq 0}$. An STL formula $\phi$ is given by the grammar $\phi::=\top\mid\pi\mid\lnot\phi\mid\phi_{1}\wedge\phi_{2}\mid\phi_{1}\mathcal{U}_{[a,b]}\phi_{2}.$ Boolean true is $\top$, and $\pi$ is a predicate of the form $\pi(s(t)):=f_{\pi}(s(t))\geq 0$ where $f_{\pi}:\mathcal{S}\to{\mathbb{R}_{e}}$ and $s(t)$ is the signal value at time instant $t$. The logical not is $\lnot$, the conjunction is $\wedge$, and $\mathcal{U}_{[a,b]}$ is the “Until” operator. Additional operators, disjunction $\vee$, Always $\square_{[a,b]}$, and Eventually $\Diamond_{[a,b]}$ can be derived from operators in the grammar [^1] Disjunction is $\phi_{1}\vee\phi_{2}=\lnot(\lnot\phi_{1}\wedge\lnot\phi_{2})$, Eventually is $\Diamond_{[a,b]}\phi=\top\mathcal{U}_{[a,b]}\phi$, and Always is $\square_{[a,b]}\phi=\lnot(\Diamond_{[a,b]}\lnot\phi)$.. Subscript ${[a,b]}$ defines the time interval. When the time interval is from 00 to $\infty$, the subscript is omitted. We will denote the set of all well-formed STL formulas with $\cal{F}$. If a signal $s$ satisfies a formula $\phi$ at time $t$, it is shown as $(s,t)\models\phi$. If it violates at $t$, it is shown as $(s,t)\not\models\phi$. The qualitative semantic of STL is defined as follows:
@@
Derived operators have following qualitative semantics:
@@
For qualitative semantics at time instant $t=0$, we omit $t$ and write $s\models\phi$. STL also has quantitative semantics to measure how well the signal models the formula. There are different quantitative semantics, also known as robustness metrics [34, 35]. In this paper we use the traditional robustness metric $\rho:\mathcal{S}\times\mathcal{F}\times\mathbb{T}\to\mathbb{R}_{e}$ from [34], defined recursively as [^2] To represent Boolean-valued quantities, we use signals $p:\mathbb{T}\to\{-\infty,+\infty\}$ and simply write $p$ as a predicate instead of $p\geq 0$. If such a signal is the $i^{th}$ coordinate of $s$ and $\pi(s(t)):=e_{i}^{T}s(t)\geq 0$ with $e_{i}$ being the $i^{th}$ natural basis vector, we have $\rho(s,\pi,t)=e_{i}^{T}s(t)=p(t)$.:
@@
The robustness value at $t=0$ is shown as $\rho(s,\phi)$. Note that for finite signals where $t_{final}<\infty$, time interval $[t+a,t+b]$ in temporal operators may exceed the time length of the signal. In this case, time interval can be taken as $[t+a,\min(t+b,t_{final})]$ assuming that $t+a\leq t_{final}$. For simplicity, we keep the semantics for infinite signals but we use STL for finite signals with necessary corrections[36]. The robustness metric $\rho$ is sound, i.e., $\rho(s,\phi,t)\geq 0\iff(s,t)\models\phi$.
@@ III-B Weighted Signal Temporal Logic (WSTL)
Note that since we have $\top$ when defining Eventually from Until, and since the WSTL robustness of $\top$ is $\infty$, we drop the set of weights $w^{2}$ in the WSTL robustness of Eventually because they do not affect the result of the min operation in the computation of the WSTL robustness of Until. That is the reason why Eventually (and hence Always) has fewer weights than Until. Moreover, the Boolean true, predicates, and the negation operator do not have associated weights, i.e., these operators have weights equal to $1$.
@@
According to Lemma 1, it is sufficient to prove that quantitative semantics in (1) is sign-consistent. Since all weights are defined positive, multiplying a robustness value with a weight does not change its sign. Therefore, for each recursive operation in the WSTL robustness calculation, the sign of the robustness value associated with this recursion step is preserved. Then, we see that $\text{sign}(\rho(s,\phi_{\rm STL},t))=\text{sign}(\tilde{r}(s,\phi,t))$ for all $(s,\phi,t)\in\mathcal{S}\times\mathcal{F}\times\mathbb{T}$ and for all non-negative weights. ∎
@@ IV Problem Statement and Solution Method
The goal of this work is to select a weight valuation $\tilde{w}$ for the parameter set $\mathcal{W}$ of the PWSTL formula $\phi_{\mathcal{W}}$. Formula $\phi_{\mathcal{W}}$ is determined according to system rules, so that it reflects safety specifications. Formally, this paper aims to solve the following problem:
@@
$\vspace{-0.2cm}r(s_{i}^{+},\phi_{\mathcal{W}=w})>r(s_{i}^{-},\phi_{\mathcal{W}=w})\quad\forall(s_{i}^{+},s_{i}^{-})\in\mathcal{P}.$ (2)
Problem 1 is a feasibility problem. In the next subsection, we reformulate it as an optimization problem to be able to handle infeasibility. Before doing so, we provide an analysis of the set of feasible weights using the syntax tree of STL formulas. An STL formula has an associated syntax tree, in which nodes represent Boolean and temporal operators, leaf nodes represent predicates, and edges represent the connection between operators and operands [37]. The syntax tree associated with the STL formula in Example 1 can be seen in Figure 1. Let the root weights of a WSTL formula be the weights associated with the weighted operator closest to the root of its syntax tree. For instance, for $\phi$ of Example 1, the root of the syntax tree is $\Diamond$ operator and the root weights are $\{w_{i}^{\Diamond}\}_{i=1}^{4}$. Consider another example with $\tilde{\phi}=\lnot\phi$, although the root operator is $\lnot$, since it is not a weighted operator, we need to look at its children until we find a weighted operator. Hence, the root weights are again $\{w_{i}^{\Diamond}\}_{i=1}^{4}$.
@@
Next, we show that the feasible weight valuations of Problem 1 is unbounded, when non-empty, due to homogeneity with respect to the root weights of the formula.
@@ IV-A An Optimization Reformulation
Problem 2 not only transforms the feasibility Problem 1 into an optimization problem, but also returns a valuation that makes maximum number of pairs correctly ordered according to Inequality (2) when Problem 1 is infeasible.
It is important to note that with Problem 2 and weights being positive, it is impossible to find weight valuations that result in a greater robustness value of a rule-violating behavior than the robustness value of a rule-satisfying one. Violating signals will always have negative robustness values. If there exists a pair in the preference dataset such that the person prefer a rule-violating behavior over a satisfying behavior, we cannot satisfy Inequality (2) for this pair, Problem 1 becomes infeasible and we will find a valuation that satisfies Inequality (2) for maximum number of pairs.
@@
We note that for certain STL formulas $\phi_{\rm STL}$, the corresponding WSTL robustness metric satisfies $r(s,\phi_{\mathcal{W}=w})\leq 0$ for all signals $s$ and all weights $w\in\mathbb{R}_{+}^{n}$. In this case, the problem of learning weights to capture preferences from only rule-satisfying signals becomes trivial since $r(s,\phi_{\mathcal{W}=w})=0$ for all $s\models\phi$. While such formulas can be uncommon in certain domains, we find them to be common in driving scenarios when the specification involves coming to a full stop or a Boolean indicator for pedestrians or traffic lights. Hence, we discuss a workaround to enable preference learning for a class of such formulas but similar workarounds can be devised for other cases too. Consider a formula of the form $\phi=\phi_{1}\wedge\square(f(s)\geq 0\;\wedge\;-f(s)\geq 0)$, second part of which essentially represents an equality $f(s)=0$. To enable preference learning in this case, we append $s$ with a new coordinate $b$ such that $b(t)=\infty$ if $f(s(t))=0$ and $b(t)=-\infty$ otherwise; and replace the formula with $\phi^{\prime}=\phi_{1}\wedge\square b$. With this transformation, if for all $t\in[0,t_{final}]$, $b(t)=\infty$, the WSTL robustness of ϕsuperscriptitalic-$\phi^{\prime}$ is determined by $\phi_{1}$; otherwise the WSTL robustness of ϕsuperscriptitalic-$\phi^{\prime}$ becomes $\infty$- ∞, which indicates a violation of $\phi$. Applying our method to ϕsuperscriptitalic-$\phi^{\prime}$ allows us to rank rule-satisfying signals.

### 12 Mar 2024 v3

@@ Abstract
This work introduces a preference learning method that ensures adherence to given specifications, with an application to autonomous vehicles. Our approach incorporates the priority ordering of Signal Temporal Logic (STL) formulas describing traffic rules into a learning framework. By leveraging Parametric Weighted Signal Temporal Logic (PWSTL), we formulate the problem of safety-guaranteed preference learning based on pairwise comparisons and propose an approach to solve this learning problem. Our approach finds a feasible valuation for the weights of the given PWSTL formula such that, with these weights, preferred signals have weighted quantitative satisfaction measures greater than their non-preferred counterparts. The feasible valuation of weights given by our approach leads to a weighted STL formula that can be used in correct-and-custom-by-construction controller synthesis. We demonstrate the performance of our method with a pilot human subject study in two different simulated driving scenarios involving a stop sign and a pedestrian crossing. Our approach yields competitive results compared to existing preference learning methods in terms of capturing preferences and notably outperforms them when safety is considered.
@@ I Introduction
Preferences are a fundamental aspect of human behavior and decision-making, and it is valuable to design autonomous systems that allow for personalization to better suit the needs and desires of users. In particular for autonomous driving, surveys have demonstrated that drivers have different comfort and performance preferences while driving in different scenarios and conditions [2, 3, 4]. Moreover, drivers tend to prefer different driving styles for autonomous vehicles than their own styles [5]. Customizing autonomous vehicles based on user preferences can increase user satisfaction with these vehicles. However, autonomous systems often require the satisfaction of a set of rules for safe operation. Relying on human preferences alone may result in unsafe behaviors. For instance, at an intersection with a stop sign, drivers may sometimes prefer a rolling stop, which is illegal, over a full stop. However, an autonomous vehicle should always stop completely at a stop sign to guarantee the safety of all agents in the environment. Preference learning algorithms for safety-critical operations must consider rule satisfaction. The main motivation for our work is the need for safe, trustworthy, and customizable autonomous vehicle algorithms.
For safety-critical applications like driving, there are three desirable properties a preference learning method should satisfy to allow safe personalization: (i) expressivity: the model should be expressive enough to capture preferences; (ii) safety: it ensures safety by preferring a rule-following behavior against a rule-violating one (even in cases where the latter is scarce in the training data); and (iii) usability in control design: the learned model should be easy to integrate into downstream correct-by-construction control synthesis tasks. In this work, we propose an integrated framework for personalization and safety to satisfy all of these properties by using Signal Temporal Logic (STL). STL is a variant of temporal logic that is tailored for reasoning about the temporal properties of time series data and is commonly used in describing correct behaviors in a variety of autonomous systems [6, 7, 8, 9, 10, 11].
@@
To evaluate the performance of our framework, we simulate two different driving scenarios: one with an autonomous vehicle navigating an intersection with a stop sign and one with an autonomous vehicle approaching a crosswalk while a pedestrian is crossing the road. We generate two sets of trajectories that comply with traffic rules for these scenarios and run a pilot human subject study with eight participants for both scenarios. Comparisons with baseline preference learning methods verify the need for safety-aware preference learning by showing that baseline methods usually lead to unsafe selections, whereas our method does not.
@@ II Literature Review
On the other hand, encoding safety rules in temporal logic is an eminent method for safety-critical applications [20, 21]. Specifications in temporal logic can be used for controller synthesis [6, 7, 21], motion planning [8, 22, 23] and learning applications [24, 9, 10, 25, 26, 11, 27, 28] in many autonomous systems. In particular, works in [24, 9, 10, 25, 29] try to infer a temporal logic formula for classification from the data. As a subset of learning applications, in robot learning, Chou et al. [26] try to learn task specifications in linear temporal logic from demonstrations, Puranic et al. [11] score demonstrations with the help of ordered specifications in the form of signal temporal logic and works in [27, 28] use temporal logic for reward shaping and reinforcement learning.
Incorporating preferences and priorities with temporal logic is studied in [12, 30, 31, 32, 1]. The work in [12] introduces a weighted variant of the STL, called Weighted Signal Temporal Logic (WSTL), in which weights reflect the order of priorities or preferences. The work in [30] defines Weighted Truncated Linear Temporal Logic. Both works assume that they have knowledge of the formula and associated weights. For the end-user, it is hard to interpret the weights and define their preferences in the temporal logic formalism, so there needs to be an intermediate step to infer the weights from the user. In [31, 32], a parametric extension of WSTL, which we call PWSTL, is used in a time series classification problem, where weights of the formula are learned using neural networks.
@@ III-A Signal Temporal Logic (STL)
STL is a temporal logic formalism used to reason about signals $s:\mathbb{T}\to\mathcal{S}$, where $\mathbb{T}$ is a time domain and $\mathcal{S}\subseteq\mathbb{R}^{m}_{e}$ is a $m$ dimensional extended real-valued signal domain [33]. We will consider $\mathbb{T}$ to be infinite $\mathbb{Z}_{\geq 0}$ or finite $[0,t_{final}]\subset\mathbb{Z}_{\geq 0}$. An STL formula $\phi$ is given by the grammar $\phi::=\top\mid\pi\mid\lnot\phi\mid\phi_{1}\wedge\phi_{2}\mid\phi_{1}\mathcal{U}_{[a,b]}\phi_{2}.$ Boolean true is $\top$, and $\pi$ is a predicate of the form $\pi(s(t)):=f_{\pi}(s(t))\geq 0$ where $f_{\pi}:\mathcal{S}\to\mathbb{R}_{e}$ and $s(t)$ is the signal value at time instant $t$. The logical not is $\lnot$, the conjunction is $\wedge$, and $\mathcal{U}_{[a,b]}$ is the “Until” operator. Additional operators, disjunction $\vee$, Always $\square_{[a,b]}$, and Eventually $\Diamond_{[a,b]}$ can be derived from operators in the grammar [^1] Disjunction is $\phi_{1}\vee\phi_{2}=\lnot(\lnot\phi_{1}\wedge\lnot\phi_{2})$, Eventually is $\Diamond_{[a,b]}\phi=\top\mathcal{U}_{[a,b]}\phi$, and Always is $\square_{[a,b]}\phi=\lnot(\Diamond_{[a,b]}\lnot\phi)$.. Subscript ${[a,b]}$ defines the time interval. When the time interval is from 00 to $\infty$, the subscript is omitted. We will denote the set of all well-formed STL formulas with $\cal{F}$. If a signal $s$ satisfies a formula $\phi$ at time $t$, it is shown as $(s,t)\models\phi$. If it violates at $t$, it is shown as $(s,t)\not\models\phi$. The qualitative semantics of STL are defined as follows:
@@
Derived operators have the following qualitative semantics:
@@
For qualitative semantics at time instant $t=0$, we omit $t$ and write $s\models\phi$. STL also has quantitative semantics to measure how well the signal models the formula. There are different quantitative semantics, also known as robustness metrics [34, 35]. In this paper, we use the traditional robustness metric $\rho:\mathcal{S}\times\mathcal{F}\times\mathbb{T}\to\mathbb{R}_{e}$ from [34], defined recursively as [^2] To represent Boolean-valued quantities, we use signals $p:\mathbb{T}\to\{-\infty,+\infty\}$ and simply write $p$ as a predicate instead of $p\geq 0$. If such a signal is the $i^{th}$ coordinate of $s$ and $\pi(s(t)):=e_{i}^{T}s(t)\geq 0$ with $e_{i}$ being the $i^{th}$ natural basis vector, we have $\rho(s,\pi,t)=e_{i}^{T}s(t)=p(t)$.:
@@
The robustness value at $t=0$ is shown as $\rho(s,\phi)$. Note that for finite signals where $t_{final}<\infty$, time interval $[t+a,t+b]$ in temporal operators may exceed the time length of the signal. In this case, time interval can be taken as $[t+a,\min(t+b,t_{final})]$ assuming that $t+a\leq t_{final}$. For simplicity, we keep the semantics for infinite signals but we use STL for finite signals with necessary corrections[36]. The robustness metric $\rho$ is sound, i.e., $\rho(s,\phi,t)>0\implies(s,t)\models\phi$ and $\rho(s,\phi,t)<0\implies(s,t)\not\models\phi$ [37].
@@ III-B Weighted Signal Temporal Logic (WSTL)
Note that since we have $\top$ when defining Eventually from Until, and since the WSTL robustness of $\top$ is $\infty$, we drop the set of weights $w^{2}$ in the WSTL robustness of Eventually because they do not affect the result of the min operation in the computation of the WSTL robustness of Until. That is the reason why Eventually (and hence Always) has fewer weights than Until. Moreover, the Boolean true, predicates, and negation operators do not have associated weights, i.e., these operators have weights equal to $1$.
@@
According to Lemma 1, it is sufficient to prove that quantitative semantics in (1) is sign-consistent. Since all weights are defined as positive, multiplying a robustness value with a weight does not change its sign. Therefore, for each recursive operation in the WSTL robustness calculation, the sign of the robustness value associated with this recursion step is preserved. Then, we see that $\text{sign}(\rho(s,\phi_{\rm STL},t))=\text{sign}(\tilde{r}(s,\phi,t))$ for all $(s,\phi,t)\in\mathcal{S}\times\mathcal{F}\times\mathbb{T}$ and for all non-negative weights. ∎
@@ IV Problem Statement and Solution Method
The goal of this work is to select a weight valuation $\tilde{w}$ for the parameter set $\mathcal{W}$ of the PWSTL formula $\phi_{\mathcal{W}}$. Formula $\phi_{\mathcal{W}}$ is determined according to system rules so that it reflects safety specifications. Formally, this paper aims to solve the following problem:
@@
$r(s_{i}^{+},\phi_{\mathcal{W}=w})>r(s_{i}^{-},\phi_{\mathcal{W}=w})\quad\forall(s_{i}^{+},s_{i}^{-})\in\mathcal{P}.$ (2)
Problem 1 is a feasibility problem. In the next subsection, we reformulate it as an optimization problem to be able to handle infeasibility. Before doing so, we provide an analysis of the set of feasible weights using the syntax tree of STL formulas. An STL formula has an associated syntax tree, in which nodes represent Boolean and temporal operators, leaf nodes represent predicates, and edges represent the connection between operators and operands [38]. The syntax tree associated with the STL formula in Example 1 is given in Figure 1. Let the root weights of a WSTL formula be the weights associated with the weighted operator closest to the root of its syntax tree. For instance, for $\phi$ of Example 1, the root of the syntax tree is $\Diamond$ operator and the root weights are $\{w_{i}^{\Diamond}\}_{i=1}^{4}$. Consider another example with $\tilde{\phi}=\lnot\phi$, although the root operator is $\lnot$ since it is not a weighted operator, we need to look at its children until we find a weighted operator. Hence, the root weights are again $\{w_{i}^{\Diamond}\}_{i=1}^{4}$.
@@
Next, we show that the feasible weight valuations of Problem 1 are unbounded, when non-empty, due to homogeneity with respect to the root weights of the formula.
@@ IV-A An Optimization Reformulation
Problem 2 not only transforms the feasibility Problem 1 into an optimization problem but also returns a valuation that makes maximum number of pairs correctly ordered according to Inequality (2) when Problem 1 is infeasible.
It is important to note that with Problem 2 and weights being positive, it is impossible to find weight valuations that result in a greater robustness value of a rule-violating behavior than the robustness value of a rule-satisfying one. Violating signals will always have negative robustness values. If there exists a pair in the preference dataset such that the person prefers a rule-violating behavior over a satisfying behavior, we cannot satisfy Inequality (2) for this pair, Problem 1 becomes infeasible and we will find a valuation that satisfies Inequality (2) for maximum number of pairs.
@@
We note that for certain STL formulas $\phi_{\rm STL}$, the corresponding WSTL robustness metric satisfies $r(s,\phi_{\mathcal{W}=w})\leq 0$ for all signals $s$ and all weights $w\in\mathbb{R}_{+}^{n}$. In this case, the problem of learning weights only from rule-satisfying signals is not meaningful since these signals will have $r(s,\phi_{\mathcal{W}=w})=0$ for any weight $w$; and satisfaction cannot be deduced when $r(s,\phi_{\mathcal{W}=w})=0$. While such formulas can be uncommon in certain domains, we find them to be common in driving scenarios when the specification involves coming to a full stop or a Boolean indicator for pedestrians or traffic lights. Hence, we discuss a workaround to enable preference learning among satisfying signals for a class of such formulas but similar workarounds can be devised for other cases too. Consider a formula of the form $\phi=\phi_{1}\wedge\square(f(s)\geq 0\;\wedge\;-f(s)\geq 0)$, second part of which essentially represents an equality $f(s)=0$. To enable preference learning in this case, we append $s$ with a new coordinate $b$ such that $b(t)=\infty$ if $f(s(t))=0$ and $b(t)=-\infty$ otherwise; and replace the formula with $\phi^{\prime}=\phi_{1}\wedge\square b$. With this transformation, if for all $t\in[0,t_{final}]$, $b(t)=\infty$, the WSTL robustness of ϕsuperscriptitalic-$\phi^{\prime}$ is determined by $\phi_{1}$; otherwise the WSTL robustness of ϕsuperscriptitalic-$\phi^{\prime}$ becomes $\infty$- ∞, which indicates a violation of $\phi$. Applying our method to ϕsuperscriptitalic-$\phi^{\prime}$ allows us to rank rule-satisfying signals.